# Time-tronics: from temporal printed circuit board to quantum computer

### v2

@@
Quantum computer. As an important example of the application of a temporal printed circuit board, we consider the case of a quantum computer. The construction of a universal quantum computer requires the preparation of numerous qubits (e.g., two-level atoms), the ability to perform all single-qubit operations, and the capability to execute two-qubit operations (e.g., controlled-Z gate operation) between any pairs of qubits [39]. The latter necessitates selective interactions between atoms, which, in turn, demand precise coherent transport of atoms and activation of interactions when they are in close proximity [49, 43, 50].
@@
The crystalline structure in time, as described in Fig. 1, consists of $s$ states ($s$ localized wave-packets). In such a structure, if we prepare $s/2$ atoms, we can define $s/2$ qubits, assigning two different wave-packets to each atom (see Fig. 3). If the atom occupies the first chosen wave-packet, the qubit will be in the state $|1\rangle$ and if it occupies the second wave-packet, it will be in the state $|0\rangle$. The initialization process of all qubits in, for example, state $|1\rangle$ is relatively straightforward. In the 1D box potential and in the presence of a static deep optical lattice, we prepare bosons in a Mott insulator state with unit filling of lattice sites [44], assuming the number of lattice sites between the walls of the box potential is $s/2$ (see Methods). Next, we turn off the interactions between atoms setting the s-wave scattering length to zero by means of a Feshbach resonance [44]. We then set the atoms moving in one direction with momentum $p$ that satisfies the $s:1$ resonance condition with the frequency of the optical lattice oscillation. To do this, we set the initially static optical lattice in motion so that in the moving frame the atoms follow the ground state of the moving potential well. This is achieved using a counter-diabatic driving process which involves a unitary transformation to the moving frame by applying a linear potential proportional to the acceleration with which the lattice is set in motion (see Supplemental Material for details). We then activate the optical lattice oscillation. This leads to a situation where $s/2$ wave-packets of the time crystalline structure that are initially moving in the same direction are occupied by individual atoms, while the other $s/2$ wave-packets initially moving in the opposite direction are unoccupied, see Fig. 3a and Methods. For each occupied wave-packet, we assign any unoccupied wave-packet to form a qubit together. The choice of which unoccupied wave-packets are assigned to which occupied wave-packets is arbitrary.
@@
If we want to perform $\sigma_{x}$ or $\sigma_{y}$ operations on a selected qubit, we have to wait for the moment when two wave-packets corresponding to the given qubit meet during the evolution along the periodic trajectory. At that moment, we activate the appropriate focused laser beam along with a second broad laser beam, allowing Bragg scattering and thus the transfer of an atom from one wave-packet moving with momentum $p$ to another wave-packet moving with momentum $-p$, or vice versa (Fig. 3a). Phase control between the laser beams allows choosing whether the transfer corresponds to $\sigma_{x}$ or $\sigma_{y}$ operations, the composition of which enables the realization of $\sigma_{z}$ operations. Very high fidelity of single-qubit operations is achieved if the full transfer of atoms between wave-packets is divided into a few encounters (see Fig. 1d and Methods). The period of the resonant trajectory can be regarded as a single clock cycle of the quantum processor. This means that a few cycles of the processor clock are needed to perform high-fidelity single-qubit operations.
So far, we have assumed that the atoms do not interact. Interaction arises if we change the internal state of the atoms. The previously described focused laser beams allow for the local activation of interactions between atoms when the wave-packets they occupy meet during the evolution along the periodic trajectory. This enables us to implement a controlled-Z gate between any pair of qubits.
@@
Let us assume that one atom occupies the wave-packet corresponding to the state $|1\rangle$ of a certain qubit, and another atom occupies the wave-packet corresponding to the state $|1\rangle$ of another qubit. Just before the moment when these two wave-packets meet, we can change the internal state of the atoms by activating the appropriate focused laser beams together with the broad laser beam, causing Raman transitions and altering the internal state of the atoms (Fig. 3b). Then, during the passage of the wave-packets, the atoms interact. Immediately after the passage, we use laser beams again to restore the atoms to their initial internal states, in which there are no interactions. If the interaction energy is properly chosen, a phase imprint of $e^{i\pi}$ occurs during the interaction. Such a phase imprint only occurs when both qubits are in the $|1\rangle$ states, thus implementing the controlled-Z gate. To achieve high-fidelity controlled-Z gates, the entire gate operation needs to be divided into a few encounters of wave-packets (Fig. 1e-1f), i.e., a few clock cycles of the quantum processor are necessary. All pairs of wave-packets meet each other during a single period, $sT$, of the resonant trajectory. Since the controlled-Z gate implementation does not alter the wave-packet populations of atoms, it can be executed simultaneously even for all possible qubit pairs, i.e., for $s/2(s/2-1)$ pairs, significantly reducing quantum computation time. The ability to perform two-qubit operations between all qubits makes a key parameter of quantum computing, i.e., the quantum volume [47], achieve very high values, especially when the Raman transfer error is reduced (Fig. 3c).
@@
Discussion and outlook
@@
I Methods
Time crystalline structure. Consider a single atom confined within a 1D box potential periodically perturbed by an oscillating optical lattice potential (Fig. 1a). Using units of length, energy, and time, denoted by $1/k$, $\hbar^{2}k^{2}/m$, and $m/\hbar k^{2}$, respectively, where $k$ is the wave number of the laser beam creating the oscillating optical lattice potential, and $m$ is the mass of the atom, the system Hamiltonian is given by
Before considering the quantum resonant behavior of the atom, let us commence with a classical description. We introduce a new pair of momentum and position variables, known as action-angle variables [29], $I=s|p|/2$ and $|\theta|=2z/s$ where $-\pi\leq\theta<\pi$. Switching to the moving frame, $\Theta=\theta-\omega t/s$, and neglecting rapidly oscillating terms, the effective Hamiltonian becomes $H_{\rm eff}=P^{2}/2+(\lambda s^{2}/32)\cos(s\Theta)$, where $P=I-\omega/s$ [29, 1]. For $s\gg 1$, $H_{\rm eff}$ describes an atom moving in a spatially periodic potential. In the quantum description, this results in solid-state-like behavior, which is manifested as the formation of energy bands and eigenstates in the form of Bloch waves [1, 30].

### v3

@@ III.1 The idea of realizing a quantum computer
The construction of a universal quantum computer requires the preparation of numerous qubits (e.g., two-level atoms), the ability to perform all single-qubit operations, and the capability to execute two-qubit operations (e.g., controlled-Z gate operation) between any pairs of qubits [39]. The latter necessitates selective interactions between atoms, which, in turn, demand precise coherent transport of atoms and activation of interactions when they are in close proximity [49, 50, 51].
@@
The crystalline structure in time, as described in Fig. 1, consists of $s$ states ($s$ localized wave-packets). In such a structure, if we prepare $s/2$ atoms, we can define $s/2$ qubits, assigning two different wave-packets to each atom (see Fig. 3). If the atom occupies the first chosen wave-packet, the qubit will be in the state $|1\rangle$ and if it occupies the second wave-packet, it will be in the state $|0\rangle$. The initialization process of all qubits in, for example, state $|1\rangle$ is relatively straightforward. In the 1D box potential and in the presence of a static deep optical lattice, we prepare bosons in a Mott insulator state with unit filling of lattice sites [44], assuming the number of lattice sites between the walls of the box potential is $s/2$. Next, we turn off the interactions between atoms setting the s-wave scattering length to zero by means of a Feshbach resonance [44]. We then set the atoms moving in one direction with momentum $p$ that satisfies the $s:1$ resonance condition with the frequency of the optical lattice oscillation. To do this, we set the initially static optical lattice in motion so that in the moving frame the atoms follow the ground state of the moving potential well. This is achieved using a counter-diabatic driving process which involves a simple unitary transformation to the moving frame by applying a linear potential proportional to the acceleration with which the lattice is set in motion (see Appendix A for details). We then activate the optical lattice oscillation. This leads to a situation where $s/2$ wave-packets of the time crystalline structure that are initially moving in the same direction are occupied by individual atoms, while the other $s/2$ wave-packets initially moving in the opposite direction are unoccupied, see Fig. 3(a). For each occupied wave-packet, we assign any unoccupied wave-packet to form a qubit together. The choice of which unoccupied wave-packets are assigned to which occupied wave-packets is arbitrary.
@@
If we want to perform $\sigma_{x}$ or $\sigma_{y}$ operations on a selected qubit, we have to wait for the moment when two wave-packets corresponding to the given qubit meet during the evolution along the periodic trajectory. At that moment, we activate the appropriate focused laser beam along with a second broad laser beam, allowing Bragg scattering and thus the transfer of an atom from one wave-packet moving with momentum $p$ to another wave-packet moving with momentum p-p, or vice versa [Fig. 3(a)]. Phase control between the laser beams allows choosing whether the transfer corresponds to $\sigma_{x}$ or $\sigma_{y}$ operations, the composition of which enables the realization of $\sigma_{z}$ operations. Very high fidelity of single-qubit operations is achieved if the full transfer of atoms between wave-packets is divided into a few encounters (see Sec. III.2 and Fig. 4). The period of the resonant trajectory can be regarded as a single clock cycle of the quantum processor. This means that a few cycles of the processor clock are needed to perform high-fidelity single-qubit operations.
@@
So far, we have assumed that the atoms do not interact. Interaction arises if we change the internal state of the atoms. The previously described focused laser beams allow for the local activation of interactions between atoms when the wave-packets they occupy meet during the evolution along the periodic trajectory. This enables us to implement a controlled-Z gate between any pair of qubits in the following way. Let us assume that one atom occupies the wave-packet corresponding to the state $|1\rangle$ of a certain qubit, and another atom occupies the wave-packet corresponding to the state $|1\rangle$ of another qubit. Just before the moment when these two wave-packets meet, we can change the internal state of the atoms by activating the appropriate focused laser beams together with the broad laser beam, causing Raman transitions and altering the internal state of the atoms [Fig. 3(b)]. Then, during the passage of the wave-packets, the atoms interact. Immediately after the passage, we use laser beams again to restore the atoms to their initial internal states, in which there are no interactions. If the interaction energy is properly chosen, a phase imprint of $e^{i\pi}$ occurs during the interaction. Such a phase imprint only occurs when both qubits are in the $|1\rangle$ states, thus implementing the controlled-Z gate. To achieve high-fidelity controlled-Z gates, the entire gate operation needs to be divided into a few encounters of wave-packets (see Sec. III.2 and Fig. 4), i.e., a few clock cycles of the quantum processor are necessary. All pairs of wave-packets meet each other during a single period, $sT$, of the resonant trajectory. Since the controlled-Z gate implementation does not alter the wave-packet populations of atoms, it can be executed simultaneously even for all possible qubit pairs, i.e., for s/2(s/21)s/2(s/2-1) pairs, significantly reducing quantum computation time. The ability to perform two-qubit operations between all qubits makes a key parameter of quantum computing, i.e., the quantum volume [52], achieve very high values, especially when the Raman transfer error is reduced [see Sec. III.2 and Fig. 4(d)].
The final stage of the quantum computation is the readout of information from the system’s final state. This can be achieved by measuring the atomic populations of the wave packets — further details are provided in Appendix B.
@@
Although our focus is primarily on bosons, the same scheme applies to fermions, which offer advantages for quantum computing due to Fermi statistics and the absence of $s$-wave scattering in a single spin state. Controlled Raman transfers enable the creation of spin mixtures, with scattering confined to selected sites. Because atoms remain in the lowest-energy state except during Raman transfers and controlled-Z gate operations, the approach is not limited by coherence times associated with spontaneous emission.
[…]
III.2 An example of a specific realization
[…]
A single focused laser beam (together with a broad beam) is sufficient to perform single-qubit operations for all qubits if the qubit states are assigned so that the wave-packets corresponding to $|0\rangle_{i}$ and $|1\rangle_{i}$ pass each other at different moments of time but at the same location for all $i$, i.e., all qubits. On the other hand, if multiple focused laser beams are available, the appropriate assignment allows for parallel single-qubit operations.
For 39K atoms, which we use as an example here, the s-wave scattering length can be set to zero if the atoms are initially prepared in the hyperfine state $|1,+1\rangle$ and the magnetic field is set to the zero crossing value of 350.5 G, in the vicinity of the Feshbach resonance at 402.5 G [54, 48]. However, if we transfer the atoms to the state $|1,-1\rangle$, which is 213 MHz above the $|1,+1\rangle$ state, using a Raman pulse, the corresponding scattering length becomes 29-29 Bohr radii, and the interaction coefficient (7) does not vanish. By varying the frequencies $\omega_{\perp}$ of the harmonic trap in the transverse directions (i.e., choosing different values of $g_{0}$), we can analyze how many meetings of the two wave-packets are needed to imprint a $\pi$ phase in the state in which each of the two atoms occupies one of the wave-packets. Simultaneously, we can determine the fidelity of the process, which is the probability that the atoms remain in their initially occupied wave-packets. The results of this analysis are presented in Figs. 4(b)-4(c). They correspond to $\omega_{\perp}/2\pi$ in the range between 50 kHz and 700 kHz.
For the parameters chosen and for the fidelity, $F_{R}=0.999933$, achieved in the Raman transfer in Ref. [50], the optimal fidelity of the controlled-Z gate is 0.993 which requires 7 cycles of the resonant trajectory [Figs. 4(b)-4(c)]. The fidelity was determined by multiplying the fidelity obtained assuming perfect Raman transfers [black curves in Figs. 4(b)-4(c)] by $F_{R}^{8n}$, where the $8n$ comes from the fact that for each of $n=7$ cycles two interactions take place where two atoms are transferred twice by Raman pulses. Higher fidelity of the gate can be achieved either by improving $F_{R}$ or by increasing the frequency of the optical lattice oscillation $\omega$. A higher $\omega$ allows for a greater oscillation amplitude $\lambda$, thereby enabling the use of stronger interactions without undesired couplings to other states of the system. Stronger interactions facilitate imprinting the $\pi$ phase onto states with high fidelity in fewer cycles of the resonant trajectory, thereby reducing the impact of imperfect Raman transfers.
A significant advantage of the proposed quantum computer is the ability to perform two-qubit gates between any pair of qubits. This is quantitatively illustrated by the quantum volume presented in Fig. 4(d). In the example we consider here and for the Raman transfer fidelity achieved in Ref. [50], $F_{R}=0.999933$, and for 49 cycle single-qubit gates [cf. Fig. 4(a)], a system of 7 qubits is capable of executing universal quantum operations. In this case, achieving the $14:1$ resonance is necessary, and for 39K atoms, any combination of the $\sigma_{x}$ and $\sigma_{y}$ operations takes 126 ms, while the controlled-Z gate operation takes 36 ms. We note that during these periods all single-qubit operations and controlled-Z gate operations, even for all possible qubit pairs, can be executed simultaneously, significantly reducing quantum computation time.
IV Summary and conclussions
@@
Acknowledgments
We acknowledge Michal Grabowski for preparing the graphics. This research was funded by the National Science Centre, Poland, Project No. 2021/42/A/ST2/00017 and the Australian Research Council Discovery Project Grants DP190100815 and DP240101590.
Appendix A Counter-diabatic driving to the initial state
This appendix details the procedure for preparing the initial state in a temporal printed circuit board. The method is broadly applicable, including to quantum computing implementations and other experiments based on the temporal circuit board.
Figure 5: Fidelity of the initial state preparation. Initially, the atom is prepared in the ground state of a static optical lattice potential. It is then accelerated to the resonant momentum $p_{\rm res}=\omega/2$ using counter-diabatic driving. The figure shows the fidelity of the desired final state as a function of the amplitude of the optical lattice oscillations. Blue dots correspond to the resonant momentum $p_{\rm res}=30$, while orange squares correspond to $p_{\rm res}=60$.
In Sec. III.1, we have outlined the procedure for preparing the initial state in the quantum computer. This process consists of several steps. Initially, $N=s/2$ atoms are prepared in a Mott insulator state within a time-independent potential composed of a box potential and $s/2$ potential wells created by a standing electromagnetic wave — each well contains exactly one localized atom. Next, we can turn off interactions between atoms by tuning an external magnetic field near a Feshbach resonance such that the scattering length becomes zero. We then set the atoms in motion with the momentum value, $p_{\rm res}$, that will satisfy the resonance condition once the oscillation of the optical lattice potential is activated. In this section, we describe how to initiate atomic motion with the correct momentum quickly and with high fidelity. The method we present is an example of counter-diabatic driving [55], which quickly enables setting the atoms in motion both in the presence and absence of interactions. For simplicity, we will assume non-interacting atoms.
$H_{0}=\frac{p^{2}}{2}+\lambda\cos^{2}(z).$ (12)
Our goal is to set the optical lattice potential in motion in such a way that the atom follows the motion of the potential well in which it is localized with the momentum $p_{\rm res}$ that matches the desired resonant value. The presence of a moving optical lattice potential ensures that the atom remains in the ground state of the well in the reference frame co-moving with the well. Setting the optical lattice in motion along a trajectory $z_{0}(t)$ can be achieved using a simple unitary transformation
$U(t)$ $=$ $\exp\left[izz_{0}^{\prime}(t)\right]\exp[-ipz_{0}(t)],$ (13)
$H_{0}=\frac{p^{2}}{2}+\lambda\cos^{2}[z-z_{0}(t)]-zz_{0}^{\prime\prime}(t),$ (14)
Initially, we have the original Hamiltonian (12), and at the final moment, we want to ensure that $z_{0}(t)=p_{\rm res}t$ and consequently $z_{0}^{\prime\prime}(t)=0$. The unitary transformation (13) provides a prescription for how we should change the Hamiltonian (14) over time to achieve the desired final result. This result can be obtained by choosing different trajectories $z_{0}(t)$ and transitioning from the initial Hamiltonian to the final one at any speed. The unitary transformation (13) guarantees that, regardless of how quickly the transition is made, the transformed solution will always satisfy the Schrödinger equation and, at the final moment when $z_{0}(t)=p_{\rm res}t$, will correspond to a wave-packet moving with the desired momentum
$U(t)\psi(z)e^{-iEt}=\psi(z-p_{\rm res}t)e^{izp_{\rm res}}e^{-iEt}.$ (15)
If we want to quickly execute the procedure for preparing a moving wave-packet, we need to use a stronger linear potential with a gradient equal to the acceleration $z_{0}^{\prime\prime}(t)$. Such a potential can be created using an inhomogeneous magnetic field. For example, for the parameters used in Sec. III.2, the preparation of the initial state for quantum computing requires that the initially static optical lattice potential is accelerated to the velocity 29 mm/s which can be achieved by introducing an appropriate frequency difference between the the counter-propergating 10.6 micron laser beams. If the preparation process is realized during 5 ms, the gradient of the linear potential needed to perform the counter-diabatic driving can be created using a magnetic field gradient of about 10 G/cm. Regardless of how quickly the excitation procedure is performed, it will always remain equally precise.
At the end of the process, we replace the moving optical lattice with an optical lattice oscillating at the resonant frequency with the amplitude $4\lambda$, thereby obtaining the desired wave-packet evolving along the resonant trajectory. Within the secular approximation (or rotating wave approximation) [26, 27], in the frame co-moving resonantly with the atom, the optical lattice appears exactly as it does in the Hamiltonian (12). In a more precise description, there are also rapidly oscillating terms, which we neglect under the secular approximation. Their presence causes the moving wave-packet obtained via the counter-diabatic driving described above to be slightly different from the exact resonant state. For the parameters chosen in this article, the fidelity is 99.8%, and it can be made arbitrarily close to 100% by selecting higher resonant momenta, i.e., higher frequencies of the optical lattice oscillations. In Fig. 5 we present the fidelity versus the amplitude of the lattice oscillations for two different oscillation frequencies $\omega$.